1. (2) Give the structure of the monomer that would be used to produce each of the following polymers. -CH2-CH-CH2-CH-CH2-CHplease answer all questions
Answer #1

Polymer Monomer CH₂-CH-CH-CH-CH2-CH- CH₂CH ? CH₂=CH Vinyl chloride (PVC) -CH-CH-CH₂=CH-CH₂-CHA CH, CHA CH=CH-ethy propylene P